# MAUGIF: Mechanism-Aware Unsupervised General Image Fusion via Dual Cross-Image Autoencoders

### v1

@@ Abstract
Image fusion aims to integrate structural and complementary information from multi-source images. However, existing fusion methods are often either highly task-specific, or general frameworks that apply uniform strategies across diverse tasks, ignoring their distinct fusion mechanisms. To address this issue, we propose a mechanism-aware unsupervised general image fusion (MAUGIF) method based on dual cross-image autoencoders. Initially, we introduce a classification of additive and multiplicative fusion according to the inherent mechanisms of different fusion tasks. Then, dual encoders map source images into a shared latent space, capturing common content while isolating modality-specific details. During the decoding phase, dual decoders act as feature injectors, selectively reintegrating the unique characteristics of each modality into the shared content for reconstruction. The modality-specific features are injected into the source image in the fusion process, generating the fused image that integrates information from both modalities. The architecture of decoders varies according to their fusion mechanisms, enhancing both performance and interpretability. Extensive experiments are conducted on diverse fusion tasks to validate the effectiveness and generalization ability of our method. The code is available at https://anonymous.4open.science/r/MAUGIF.
@@ 1 Introduction
Image fusion aims to integrate complementary information from multiple source images into a single output that is more informative, visually interpretable, and suitable for downstream tasks such as object detection [He2023], segmentation [Wang2025DiFusionSegDS], and remote sensing analysis [Hong2018]. General image fusion refers to designing a unified model or framework that can be applied to various types of image fusion tasks. These tasks include hyperspectral-multispectral image fusion (HMF), visible-infrared image fusion (VIF), multi-focus image fusion (MFF), and medical image fusion (MEF). The HMF task seeks to fuse a low-spatial-resolution hyperspectral image (HSI) with a high-spatial-resolution multispectral image (MSI) to produce a high-spatial-resolution HSI [Hysure]. The VIF integrates thermal radiation data from infrared sensors with rich texture and color information from visible cameras. This results in the fused images that simultaneously highlight salient objects and retain visual fidelity [Ma2021GANMcCAG]. The MFF task aims at generating an all-in-focus image from a series of partially focused images [Quan2025]. The purpose of MEF is to display the fine structure and functional metabolism of the lesion simultaneously, thereby assisting doctors in making more accurate diagnoses and assessment [HE2025102666].
@@
Existing image fusion methods can be broadly categorized into task-specific and general-purpose. Task-specific methods are specially designed for a particular type of image fusion task. These methods typically leverage prior image information such as sparsity [OTD] and structural sparsity [S4], or deep image features [ZSL] to enhance the fusion performance [Zhang2023]. However, these methods often exhibit limited generalization capability across different fusion tasks. In contrast, general image fusion methods aim to adopt a uniform fusion strategy across tasks [Xu2020U2FusionAU]. These methods generally use general features including frequency coefficients [Wang2024a] and deep features [Liang2024FusionFD] to optimize the model. Nevertheless, most existing methods fail to fully account for the underlying mechanisms of different tasks, thereby limiting their adaptability and interpretability [Zhu2024TaskCustomizedMO].
@@
  • A mechanism-aware taxonomy of fusion tasks. By proposing a classification of fusion into additive and multiplicative types, we can gain insight into how the modality-specific features are integrated. This not only guides the architectural design of the network, but also enhances model generalization and interpretability.
@@
Figure 2: Visual illustration of the fusion process for MFF.
@@ 2.1 Task-specific image fusion methods
Task-specific fusion methods are designed for a particular fusion task, leveraging the inherent properties of image and priors to enhance performance. Model-based fusion methods typically rely on mathematical modeling and optimization to exploit the priors of images [NSSR, OTD]. For example, these methods often leverage various matrix or tensor decompositions to capture and exploit the intrinsic structure of image data [CP1, CSTF]. In addition, they can incorporate image priors via designed regularizers, such as low-rankness [Hysure], sparsity [WANG2014477], smoothness [Ma2016], and non-local self-similarity [LTMR]. Deep learning-based methods capture high-level semantic content by learning hierarchical representations directly from data. Numerous methods employ convolutional neural networks (CNNs) to extract such features from source images [ZSL, Li2019]. Subsequently, generative adversarial networks (GANs) [Ma2019FusionGANAG, Zhang2021a] and vision Transformers [Shao2024, Liu2024] have been introduced to better preserve structural details and global contextual information. More recently, diffusion-based fusion methods have also attracted significant attention [Yue2023, Wu2025]. Furthermore, high-level vision tasks including object detection and semantic segmentation, have been incorporated to guide the fusion process [Tang2022ImageFI, Wang2025DiFusionSegDS]. These methods often deliver optimal performance on specific fusion tasks but exhibit limited generalizability.
@@ 2.2 General image fusion methods
General image fusion aims to handle multiple fusion tasks within a unified framework [Ma2022SwinFusionCL]. Recently, Zhang et al. [Zhang2020a] proposed a supervised training framework for MFF, which can be generalized to other tasks by adjusting fusion conditions. Ma et al. [Ma2022] employed a unified network architecture and loss function, but trained separate models for each individual task. Wang et al. [Wang2024a] proposed a semi-supervised learning approach, which decomposes the source images into features of different frequencies and adaptively merges the features. The work in [Zhu2024] accommodated diverse fusion tasks by dynamically customizing a mixture of task-specific adapters, preserving task individuality while leveraging cross-task commonalities.
@@ 3.1 Mechanism-Aware Image Fusion
To capture the diverse fusion mechanisms of information integration, we classify image fusion into additive and multiplicative types. The classification depends on whether the injected modality-specific features are added to the target or modulate it. The specific definitions are as follows.
@@
where $\psi$ is a nonexpansive projection, i.e., $\|\psi(x)-\psi(y)\|\leq\|x-y\|$, $\Delta_{x}:=\textbf{X}-\textbf{C}$ and $\Delta_{y}:=\textbf{Y}-\textbf{C}$ are the modality-specific features unique to X and Y, respectively, and C is the common content of the same size as X and Y.
@@
(The choice of $\psi$) The design of $\psi$ is adapted to the specific fusion task. For MFF where the modality-specific information of images is all needed, $\psi$ can be set to the identity mapping, i.e. $\psi(x)=x$, to preserve all fine details. For VIF, a hard-thresholding:
@@
can be selected to suppress weak or spurious responses and enhance salient thermal targets, where $\sigma\in[0,~0.4]$.
@@ 3.2 Dual cross-image autoencoders framework
1) To extract the common components of source images, we employ two encoders, denoted as $E_{x}(\cdot)$ and $E_{y}(\cdot)$, to encode X and Y respectively, such that $E_{x}(\textbf{X})\approx E_{y}(\textbf{Y})$. The loss function for the encoders can be formulated as:
@@
2) Two decoders, denoted as $D_{x}(\cdot)$ and $D_{y}(\cdot)$, are employed to decode the learned shared content. The objective is to reconstruct the source images X and Y from the common components, i.e.,
@@
The decoder acts as a ‘feature injector’, which injects the source image’s modality-specific features into the shared content, thereby reconstructing the source images. The loss function for the decoders can be defined as:
@@
3) Finally, we perform image fusion by injecting the modality-specific feature of image X into image Y, or vice versa. Therefore, the fused image F can be formulated as:
@@
In this process, the decoder serves directly as a fusion operator, shifting the fusion process from conventional rule-based design to intrinsic feature injection.
Additive fusion and multiplicative fusion correspond to different decoder architectures, as shown in Figure 3 (e). In additive fusion, the decoder architecture follows a residual paradigm, where the modality-specific features of X are injected as an additive manner, i.e.,
@@
Under DCIAE framework, image fusion is no longer a ‘black-box’ process. We can clearly observe the modality-specific features contributed by the source images during the fusion process. Moreover, DCIAE framework is few shot, requiring only one image pair for model training. In addition, the fusion process involves just a single decoding step, making it highly efficient.
@@
  • $\mathcal{C}^{*}$: the ground-truth shared content between X and Y.
@@ 4.1.1 Implementation details
All experiments are conducted on Ubuntu 20.04 with the PyTorch 2.2.0. The hardware configuration includes an Intel Core i7-13700KF CPU with a clock speed of 3.4GHz, 32GB of RAM, and an NVIDIA GeForce RTX 4090 GPU with 24GB of VRAM. Throughout the training process, the AdamW [Loshchilov2017DecoupledWD] optimizer is employed, along with a linear learning rate schedule that incorporates a warm-up phase. The maximum learning rate is set to $1\times 10^{-4}$, the batch size is set to 32, and the training runs for 50 epochs.
@@ 4.1.2 Dataset
CAVE1 and Harvard Dataset [data11] for HMF: For the CAVE dataset, we selected ‘Toys’ and ‘Peppers’, whose spatial dimensions were downsampled by different scale factors (sf). We simulated the MSIs with three bands by the Nikon D700 camera2. For Harvard dataset, the ‘img3’ and ‘imgc6’ images are selected. Each contains a 31-band HSI, which serves as the ground truth. Then, the low-resolution HSIs are generated through downsampling the original HSI at different scale factors. We divide the spectral bands of HSI into three parts and generate a 3-band MSI by averaging the bands. We corrupt the HSI and MSI with Gaussian noise to obtain SNRs of 35 dB and 40 dB, respectively.
MSRS Dataset [Tang2022ImageFI] for VIF: This dataset is a benchmark for infrared-visible image fusion and segmentation. It comprises 1,444 well-registered image pairs. Following the official protocol, we adopt the original split with 1,083 pairs for training and 361 pairs for testing. All images have a resolution of 640×480 and are accompanied by pixel-level semantic segmentation annotations.
@@
Harvard Dataset3 for MEF: This dataset is a well-known medical imaging benchmark for multimodal fusion. It provides a large number of CT–MRI, PET–MRI, and SPECT–MRI image pairs, which are widely used for training and evaluating medical image fusion models.
@@ 4.2 Comparison on HMF task
Table 1: The evaluation metrics of the fused images for HMF.
@@
Table 1 presents the fusion performance of the compared methods on four HSI datasets. From the table, the CNMF and HySure exhibit relatively weaker performance, particularly under large scale factor cases. In this cases, the fusion model becomes severely ill-posed, which limits their performance. TriDLR reduces the number of parameters through tensor triple decomposition. It adopts a strategy of estimating the degradation operator and performing fusion simultaneously, effectively avoiding error accumulation and achieving superior performance. The ZSL achieves a relatively accurate estimation of the degradation operators, which results in less error accumulation during its subsequent fusion process and leads to competitive fusion performance. However, when the scale factor (sf) is large, the accuracy of the degradation operator estimated by ZSL decreases, leading to a significant drop in fusion performance. In contrast, the proposed MAUGIF method exhibits greater robustness to large sf, as the proposed method can perform spectral super-resolution on the MSI, which is less sensitive to spatial downsampling. Overall, MAUGIF achieves the optimal performance in four datasets, which demonstrates its effectiveness in HMF.
@@
Figure 5 presents the fused images and error maps on the ‘Toys’ dataset for visual comparison. As shown in the figure, the fused images of HySure exhibit noticeable color distortions, which indicates a significant estimation error in the spectral degradation operator. As can be seen from the error maps, ZSL achieves competitive results. However, MAUGIF exhibits the fewest artifacts and color distortion, achieving the best visual reconstruction quality.
@@ 4.3 Comparison on VIF task
To evaluate the performance of the MAUGIF on visible-infrared image fusion task, we conduct experiments on MSRS [Tang2022ImageFI] dataset, and compare the MAUGIF with several state-of-the-art (SOTA) fusion methods. The quantitative results presented in Figure 2 demonstrate that MAUGIF achieves five top rankings and one second-place ranking among all competing methods. Specifically, it records an MI of 5.638, an SCD of 1.731, a VIF of 0.949, a $Q^{AB/F}$ score of 0.652, and an SSIM of 0.482. These results indicate that MAUGIF effectively preserves mutual information, structural similarity, and visual information fidelity, suggesting superior retention of texture, structure, and perceptual quality.
@@
To better illustrate the fusion performance of MAUGIF, we provide a qualitative comparison, as shown in Figure 6. We can observe that our method preserves fine-grained texture details more effectively than compared approaches. Specifically, in the regions highlighted by the red and green boxes, the text remains sharp, structurally continuous, and clearly recognizable in our results. In contrast, other methods tend to introduce noise interference from the infrared modality, resulting in blurred textures, weakened edge contours, and noticeable degradation of structural fidelity. These visual comparisons further verify the robustness of our fusion strategy in maintaining both clarity and semantic consistency.
In summary, both quantitative and qualitative results demonstrate that MAUGIF preserves complementary information from the source images while maintaining visual fidelity. The fused images retain key details from both visible and infrared modalities, ensuring high visual quality and enhancing the overall integrity of the fusion process.

### v2

@@ Abstract
Image fusion aims to integrate structural and complementary information from multi-source images. However, existing fusion methods are often either highly task-specific, or general frameworks that apply uniform strategies across diverse tasks, ignoring their distinct fusion mechanisms. To address this issue, we propose a mechanism-aware unsupervised general image fusion (MAUGIF) method based on dual cross-image autoencoders. Initially, we introduce a classification of additive and multiplicative fusion according to the inherent mechanisms of different fusion tasks. Then, dual encoders map source images into a shared latent space, capturing common content while isolating modality-specific details. During the decoding phase, dual decoders act as feature injectors, selectively reintegrating the unique characteristics of each modality into the shared content for reconstruction. The modality-specific features are injected into the source image in the fusion process, generating the fused image that integrates information from both modalities. The architecture of decoders varies according to their fusion mechanisms, enhancing both performance and interpretability. Extensive experiments are conducted on diverse fusion tasks to validate the effectiveness and generalization ability of our method. The code is available at https://github.com/warren-wzw/MAUGIF.git.
@@ 1 Introduction
Image fusion aims to integrate complementary information from multiple source images into a single output that is more informative, visually interpretable, and suitable for downstream tasks such as object detection [He2023], segmentation [Wang2025DiFusionSegDS], and remote sensing analysis [Hong2018]. General image fusion refers to designing a unified model or framework that can be applied to various types of image fusion tasks. These tasks include hyperspectral-multispectral image fusion (HMF), visible-infrared image fusion (VIF), multi-focus image fusion (MFF), and medical image fusion (MEF). The HMF task seeks to fuse a low-spatial-resolution hyperspectral image (HSI) with a high-spatial-resolution multispectral image (MSI) to produce a high-spatial-resolution HSI [Hysure]. The VIF integrates thermal radiation data from infrared sensors with rich texture and color information from visible cameras. This results in fused images that simultaneously highlight salient objects and retain visual fidelity [Ma2021GANMcCAG]. The MFF task aims at generating an all-in-focus image from a series of partially focused images [Quan2025]. The purpose of MEF is to display the fine structure and functional metabolism of the lesion simultaneously, thereby assisting doctors in making more accurate diagnoses and assessment [HE2025102666].
@@
Existing image fusion methods can be broadly categorized into task-specific and general-purpose. Task-specific methods are specially designed for a particular type of image fusion task. These methods typically leverage prior image information such as sparsity [OTD] and structural sparsity [S4], or deep image features [ZSL] to enhance fusion performance [Zhang2023]. However, these methods often exhibit limited generalization capability across different fusion tasks. In contrast, general image fusion methods aim to adopt a uniform fusion strategy across tasks [Xu2020U2FusionAU]. These methods generally use general features including frequency coefficients [Wang2024a] and deep features [Liang2024FusionFD] to optimize the model. Nevertheless, most existing methods fail to fully account for the underlying mechanisms of different tasks, limiting their adaptability and interpretability [Zhu2024TaskCustomizedMO].
@@
  • A mechanism-aware taxonomy of fusion tasks. By proposing a classification of fusion into additive and multiplicative types, we can gain insight into how the modality-specific features are integrated. This not only guides the architectural design of the network, but also enhances the model’s generalization and interpretability.
@@
Figure 2: Visual illustration of the fusion process for MFF task.
@@ 2.1 Task-specific image fusion methods
Task-specific fusion methods are designed for a particular fusion task, leveraging the inherent properties of input images and prior knowledge to enhance performance. Model-based fusion methods typically rely on mathematical modeling and optimization to exploit image priors [NSSR, OTD]. For example, these methods often leverage various matrix or tensor decompositions to capture and exploit the intrinsic structure of image data [CP1, CSTF]. In addition, they can incorporate image priors using handcrafted regularizers, such as low-rankness [Hysure], sparsity [WANG2014477], smoothness [Ma2016], and nonlocal self-similarity [LTMR]. Deep learning-based methods capture high-level semantic content by learning hierarchical representations directly from images. Numerous methods employ convolutional neural networks (CNN) to extract such features from source images [ZSL, Li2019]. Subsequently, generative adversarial networks (GANs) [Ma2019FusionGANAG, Zhang2021a] and vision Transformers [Shao2024, Liu2024] have been introduced to better preserve structural details and global contextual information. Diffusion-based fusion methods have also attracted significant attention [Yue2023, Wu2025]. Furthermore, high-level vision tasks including object detection and semantic segmentation, have been incorporated to guide the fusion process [Tang2022ImageFI, Wang2025DiFusionSegDS]. These methods often deliver high performance on specific fusion tasks but exhibit limited generalizability.
@@ 2.2 General image fusion methods
General image fusion aims to handle multiple fusion tasks within a unified framework [Ma2022SwinFusionCL]. Recently, Zhang et al. [Zhang2020a] proposed a supervised training framework for MFF task, which can be generalized to other tasks by adjusting fusion conditions. Ma et al. [Ma2022] employed a unified network architecture and loss function, but trained separate models for each individual task. Wang et al. [Wang2024a] proposed a semi-supervised learning approach, which decomposes the source images into features of different frequencies and adaptively merges the features. The work in [Zhu2024] accommodated diverse fusion tasks by dynamically customizing a mixture of task-specific adapters, preserving task individuality while leveraging cross-task commonalities.
@@ 3.1 Mechanism-Aware Image Fusion
To capture the diverse fusion mechanisms of information integration, we classify image fusion into additive and multiplicative types. The classification depends on whether modality-specific features are added to the target or modulate it. The specific definitions are as follows.
@@
where $\psi$ is a non-expansive projection, i.e., $\|\psi(x)-\psi(y)\|\leq\|x-y\|$, $\Delta_{x}:=\textbf{X}-\textbf{C}$ and $\Delta_{y}:=\textbf{Y}-\textbf{C}$ are modality-specific features unique to X and Y, respectively, and C is the common content of the same size as X and Y.
@@
(The choice of $\psi$) The design of $\psi$ is adapted to the specific fusion task. For MFF where the modality-specific information of images is all needed, $\psi$ can be set to the identity mapping, i.e., $\psi(x)=x$, to preserve all fine details. For VIF, a hard-thresholding:
@@
can be employed to suppress weak or spurious responses and enhance salient thermal targets, where $\sigma\in[0,~0.4]$.
@@ 3.2 Dual cross-image autoencoders framework
1) To extract the common components of source images, we employ two encoders, denoted as $E_{x}(\cdot)$ and $E_{y}(\cdot)$, to encode X and Y, respectively, such that $E_{x}(\textbf{X})\approx E_{y}(\textbf{Y})$. The loss function for the encoders can be formulated as:
@@
2) Two decoders, denoted by $D_{x}(\cdot)$ and $D_{y}(\cdot)$, are employed to decode the learned shared content. The objective is to reconstruct the source images X and Y from the common components, i.e.,
@@
The decoder acts as a feature injector, which injects the source image’s modality-specific features into the shared content, thereby reconstructing the source images. The loss function for the decoders can be defined as:
@@
3) Finally, we perform image fusion by injecting the modality-specific features of image X into image Y, or vice versa. Therefore, the fused image F can be formulated as:
@@
In this process, the decoder serves directly as a fusion operator, shifting the fusion process from conventional rule-based design to the injection of modality specific feature.
Additive fusion and multiplicative fusion correspond to different decoder architectures, as shown in Figure 3 (e). In additive fusion, the decoder architecture follows a residual paradigm, where the modality-specific feature of X is injected as an additive manner, i.e.,
@@
In DCIAE framework, image fusion is no longer a ‘black-box’ process. We can clearly observe the modality-specific features contributed by the source images during the fusion process. Moreover, DCIAE framework is few shot, requiring only one image pair for model training. In addition, the fusion process involves just a single decoding step, making it highly efficient.
@@
  • $\mathcal{C}^{*}$: the ground truth shared content between X and Y.
@@ 4.1.1 Implementation details
All experiments are conducted on Ubuntu 20.04 with PyTorch 2.2.0. The hardware configuration includes an Intel Core i7-13700KF CPU with a clock speed of 3.4GHz, 32GB of RAM, and an NVIDIA GeForce RTX 4090 GPU with 24GB of VRAM. Throughout the training process, the AdamW [Loshchilov2017DecoupledWD] optimizer is employed, along with a linear learning rate schedule that incorporates a warm-up phase. The maximum learning rate is set to $1\times 10^{-4}$, the batch size is set to 32, and the training runs for 50 epochs.
@@ 4.1.2 Dataset
CAVE1 and Harvard Dataset [data11] for HMF: For the CAVE dataset, we selected ‘Toys’ and ‘Peppers’, whose spatial dimensions were downsampled by different scale factors (sf). We simulated MSIs with three bands by the Nikon D700 camera2. For the Harvard dataset, images ‘img3’ and ‘imgc6’ are selected. Each contains a 31-band HSI, which serves as the ground truth. The low-resolution HSIs are generated through downsampling the original HSI at different scale factors. We divide the spectral bands of HSI into three parts and generate a 3-band MSI by averaging the bands. We corrupt HSI and MSI with Gaussian noise to obtain SNRs of 35 dB and 40 dB, respectively.
MSRS Dataset [Tang2022ImageFI] for VIF: This dataset is a benchmark for the fusion and segmentation of infrared-visible images. It comprises 1,444 well-registered image pairs. Following the official protocol, we adopt the original split with 1,083 pairs for training and 361 pairs for testing. All images have a resolution of 640×480 and are accompanied by pixel-level semantic segmentation annotations.
@@
Harvard Dataset3 for MEF: This dataset is a well-known medical imaging benchmark for multimodal fusion. It provides a large number of CT–MRI, PET–MRI, and SPECT–MRI image pairs, which are widely used to train and evaluate medical image fusion models.
@@ 4.2 Comparison on HMF task
Table 1: The evaluation metrics of the fused images for HMF task.
@@
Table 1 presents the fusion performance of the compared methods in four HSI datasets. From the table, the CNMF and HySure exhibit relatively weaker performance, particularly in large scale factor cases. In these cases, the fusion model becomes severely ill-posed, which limits their performance. TriDLR reduces the number of parameters through tensor triple decomposition. It adopts a strategy of estimating the degradation operator and performing fusion simultaneously, effectively avoiding error accumulation and achieving superior performance. The ZSL achieves a relatively accurate estimation of the degradation operators, which results in less error accumulation during its subsequent fusion process and leads to competitive fusion performance. However, when the scale factor (sf) is large, the accuracy of the degradation operator estimated by ZSL decreases, leading to a significant drop in fusion performance. In contrast, the proposed MAUGIF method exhibits greater robustness to large sf, as the proposed method can perform spectral super-resolution on the MSI, which is less sensitive to spatial downsampling. Overall, MAUGIF achieves optimal performance in four datasets, which demonstrates its effectiveness in HMF task.
@@
Figure 5 presents the fused images and error maps in the ‘Toys’ dataset for visual comparison. As shown in the figure, the fused images of HySure exhibit noticeable color distortions, which indicates a significant estimation error in the spectral degradation operator. As can be seen from the error maps, ZSL achieves competitive results. However, MAUGIF exhibits the fewest artifacts and color distortion, achieving the best visual reconstruction quality.
@@ 4.3 Comparison on VIF task
To evaluate the performance of MAUGIF on visible-infrared image fusion task, we conduct experiments on MSRS [Tang2022ImageFI] dataset, and compare the MAUGIF with several state-of-the-art (SOTA) fusion methods. The quantitative results presented in Table 2 demonstrate that MAUGIF achieves five top rankings and one second-place ranking among all competing methods. Specifically, it records an MI of 5.638, an SCD of 1.731, a VIF of 0.949, a $Q^{AB/F}$ score of 0.652, and an SSIM of 0.482. These results indicate that MAUGIF effectively preserves mutual information, structural similarity, and visual information fidelity, suggesting superior retention of texture, structure, and perceptual quality.
@@
To better illustrate the fusion performance of MAUGIF, we provide a qualitative comparison, as shown in Figure 6. We can observe that our method preserves fine-grained texture details more effectively than compared approaches. Specifically, in the regions highlighted by the red and green boxes, the text remains sharp, structurally continuous, and clearly recognizable in our results. In contrast, other methods tend to introduce noise interference from the infrared modality, resulting in blurred textures, weakened edge contours, and noticeable degradation of structural fidelity. These visual comparisons further verify the robustness of our fusion strategy in maintaining both clarity and semantic consistency. In summary, both quantitative and qualitative results demonstrate that MAUGIF preserves complementary information from the source images while maintaining visual fidelity.